OpenStudy (anonymous):

Mixture problem: A tank contains 1000 L of solution consisting of 100kg of salt dissolved in water. Pure water is pumped into the tank at a rate of 5 L/s, and the mixture, kept uniform by stirring, is pumped out at the same rate. How long until only 10 kg of salt remains in the tank?

OpenStudy (anonymous):

So, I think I got this, but not really sure cus it seemed like it didn't really involve any differential equation...but, we would be looking time t when x(t) = 10 right? Then, \[\frac{dy}{dx} = (rate in) - (rate out)\] where the rate in would be \[\frac{0kg}{second} = 0\]and rate out would be \[\frac{5L}{second}*\frac{x(t)}{100-t} = \frac{5L}{second}*\frac{10}{100-t}=\frac{50}{100-t}\]

OpenStudy (anonymous):

so, the equation would become \[\frac{dx}{dt} = \frac{50}{100-t}\] and you would just integrate both sides and solve for t right?

OpenStudy (anonymous):

(sorry, bottom should be 100+t not minus...)

OpenStudy (anonymous):

\[\begin{align*}\frac{dy}{dt}&=\text{(rate in)}-\text{(rate out)}\\ &=\text{(volume in)}\times\text{(concentration in)}-\text{(volume out)}\times\text{(conc. out)}\\ &=5\times0-5\times\frac{y}{1000+(5-5)t}\\ &=-\frac{5y}{1000+t} \end{align*}\] This is definitely a differential equation... A DE is any equation that contains a derivative. Do you need help solving it?